OpenStudy (anonymous):

Write in terms of sine and cosine and simplify the expression. sin^4 A − cos^4 A cos A − sin A

OpenStudy (anonymous):

-cos2A

OpenStudy (anonymous):

wrong answer

OpenStudy (anonymous):

is it (sina+cosA)(sinA-cosA)

OpenStudy (dumbcow):

\[\sin^{4} -\cos^{4} =(\sin^{2}+\cos^{2})(\sin^{2}-\cos^{2}) = \sin^{2}-\cos^{2} = -\cos 2A\]